OpenStudy (anonymous):

Use logarithmic differentiation to find the derivative of the function. y = x^2x

terenzreignz (terenzreignz):

Hi :) When you get both bases and exponents as functions of x, usually, logarithmic differentiation is your only feasible choice: \[\Large y = x^{2x}\] Start by taking the ln of both sides... \[\Large \ln(y) = \ln\left(x^{2x}\right)\] can you take it from here?

OpenStudy (anonymous):

Hi, I am confused about how logarithms work. i understand how to normally find derivative, but not using ln functions

terenzreignz (terenzreignz):

Well, there are two reasons to use the logarithmic functions: One, it turns products into sums, which are much easier to differentiate. (generally) and two, the log (or ln) itself has a rather simple derivative: \[\Large \frac{d}{dt}\ln(t) = \frac1t\]

terenzreignz (terenzreignz):

Have you ever heard of implicit differentiation?

OpenStudy (anonymous):

Briefly...

terenzreignz (terenzreignz):

Well, thankfully, this particular case is not a very complicated case of implicit differentiation. In particular, with THIS in mind: \[\Large \frac{d}{dt}\ln(t) = \frac1t\] what is the derivative of the left side, with respect to x? \[\Large \color{blue}{\ln(y) }= \ln\left(x^{2x}\right)\]

OpenStudy (anonymous):

would it be 1/y

terenzreignz (terenzreignz):

1/y, but since we're differentiating with respect to x (with implicit differentiation in mind) it better be \[\Large \frac1y \cdot \frac{dy}{dx}\]or\[\Large \frac1y\cdot y'\]depending on how you were taught to differentiate implicitly.... So, catch me so far? ^_^

OpenStudy (anonymous):

yes making sense

terenzreignz (terenzreignz):

Okay, so much for the left-hand side. Let's do the right, and differentiate it. \[\Large \ln\left(x^{2x}\right)\] Normally, when both the base and exponent are functions of x, it's differentiating is not straightforward, if not downright impossible... However, since we have a log (ln), we can use one of its more beloved properties, THIS one... \[\Large \ln b^p = p\ln b\]

terenzreignz (terenzreignz):

IE, the exponent may 'go down' and simply be multiplied to the log... with that in mind, how would you 'deal' with \[\Large \ln\left(x^{2x}\right)\]

OpenStudy (anonymous):

2xlnx?

terenzreignz (terenzreignz):

You still here? Sorry, OS gave up on me T.T You're right with 2x ln (x) by the way... \[\Large 2x \ln(x)\] Now, you can use the product rule to differentiate this, and also, once again recall that \[\Large \frac{d}{dt}\ln(t) = \frac1t\]